OpenStudy (sugarmochi):

Find a,b if the line ax + by + 3 = 0 passes through (0,1) and is parallel to 2x-3y+5=0

OpenStudy (triciaal):

Parallel lines have the same slope. in the form y = m x + b the slope is m any 2 points on a given line will have the same slope

OpenStudy (mrnood):

2x-3y+5=0 can you arrange this into the form y = mx+b if so it will tell you the slope of the line that you need to find (as @triciaal has explained above)

OpenStudy (mrnood):

hint - first add 3y to each side of the equation..

OpenStudy (sshayer):

eq. of any line parallel to px+qy+r=0 is px+qy+d=0 where d is a constant.

OpenStudy (welshfella):

using sshayer's equations- 2x - 3y + 5 = 0 so the equation parallel to it will be 2x - 3y + d = 0 where d is a constant to be found

OpenStudy (welshfella):

Now if we plug in the given point (0,1) [ x = 0 and y = 1] into the second equation we can find the value of d.

OpenStudy (sugarmochi):

sorry for the late answer! I wasnt able to check my question untill now... Uhm, would d= -3?

OpenStudy (sugarmochi):

@welshfella (is this how i tag? ohmygod im sorry)

OpenStudy (sshayer):

d=3

OpenStudy (sugarmochi):

ohh- my bad